OpenStudy (anonymous):

MEDAL FOR BEST ANSWER Find the area of a sector with a central angle of 110° and a diameter of 5.2 cm. Round to the nearest tenth.

OpenStudy (wolf1728):

radius = 2.6 cm. central angle = 110 which is 0.3055555556 of a circle (110/360) so sector area = PI*r² * (110/360)

OpenStudy (anonymous):

where does the Pl come from and the radi is half of 5.2 right?

OpenStudy (wolf1728):

yes radius = 1/2 diameter

OpenStudy (wolf1728):

Are you familiar with PI at all?

OpenStudy (anonymous):

no... teach didnt teach that

OpenStudy (wolf1728):

PI is involved with all kinds of formulas dealing with circles and spheres. For example if a circle has a diameter of 10, what is the circumference?

OpenStudy (wolf1728):

PI = 3.14159265

OpenStudy (anonymous):

oooh pi i thought that was some technicality derp...

OpenStudy (wolf1728):

so circumference = PI * diameter = 2 * radius * PI

OpenStudy (anonymous):

so i got 2.4958 but thats not one my answer choices

OpenStudy (anonymous):

the answer should be 6.5

OpenStudy (wolf1728):

So I guess you are talking about PI * r² for the sector, right?

OpenStudy (anonymous):

i think so

OpenStudy (wolf1728):

So, can you calculate PI*r² * (110/360)?

OpenStudy (anonymous):

oh i got it ! I made an error. Forgot to square the radi

OpenStudy (wolf1728):

Okay then you've got it.